OpenStudy (anonymous):

y-7+7=6/5(x-3)+7 answer is y=6/5x+17/5....how did they get 17/5 ? 12:59 am

OpenStudy (anonymous):

Subtract 7 from both sides to get, \[y-7=\frac{6}{5}(x-3)\]Expand the right-hand side:\[y-7=\frac{6}{5}x-\frac{18}{5}\]When you add 7 to both sides you end up with,\[y=\frac{6}{5}x-\frac{18}{5}+7\]and you get your answer when you add the constants together:\[y=\frac{6}{5}x+\frac{17}{5}\]Hope this helps :) Feel free to 'fan' me!